find three consecutive odd positive integers such that 5 times the third integer equals three times the first integer plus seven more than the second integer

Found 3 solutions by CubeyThePenguin, MathTherapy, ikleyn:
Answer by CubeyThePenguin(3113)

consecutive odd integers: (x-2), x, (x+2)

5(x+2) = 3(x-2) + 7 + x
5x + 10 = 4x + 3
x = -7

The integers are -9, -7, and -5.

Answer by MathTherapy(10552)

find three consecutive odd positive integers such that 5 times the third integer equals three times the first integer plus seven more than the second integer
That person is WRONG, once AGAIN!! 


Answer by ikleyn(52787)

.

Let consecutive odd integers be (x-2), x, (x+2)

5(x+2) = 3(x-2) + 7 + x
5x + 10 = 4x + 1
x = -9

The integers are -11, -9, and -7, in ascending order.             ANSWER